OpenStudy (anonymous):

Choose the correct procedure used to graph the equation 2x + 4y = 24 using the intercepts method. Solve the equation for y to get a y-intercept of (0, 6) and a slope of -one half. Plot the y-intercept, count the slope to get another point and draw the line that passes through the two points. Let x equal zero and solve for y to find a y-intercept of (0, 6). Let y equal zero and solve for x to find an x-intercept of (12, 0). Plot the two points and draw a line passing through the two points. Choose two different values; plug them into the equation to find the corresponding y-values. Plo

OpenStudy (anonymous):

Choose two different values; plug them into the equation to find the corresponding y-values. Plot the two points and draw the line that passes through the two points. Solve the equation for y to find an intercept of (6, 0); solve the equation for x to find a y-intercept of (0, 12). Plot the two points and draw a line passing through the two points. The last two ^^^

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

you need two points to plot any line

jimthompson5910 (jim_thompson5910):

in this case, you just need the x and y intercepts

OpenStudy (anonymous):

y = 6 but i don't know what x is

jimthompson5910 (jim_thompson5910):

to find the x-intercept, plug in y = 0 and solve for x

OpenStudy (anonymous):

i'm very confused. i'm horrible at this.

jimthompson5910 (jim_thompson5910):

2x + 4y = 24 2x + 4(0) = 24 ... plug in y = 0 2x + 0 = 24 2x = 24 x = ??

OpenStudy (anonymous):

x is 12.

jimthompson5910 (jim_thompson5910):

so that's the x-intercept

jimthompson5910 (jim_thompson5910):

now that you know the x and y intercept, you have your two points needed to draw the line

OpenStudy (anonymous):

So, the very last one, right?

jimthompson5910 (jim_thompson5910):

no it's Let x equal zero and solve for y to find a y-intercept of (0, 6). Let y equal zero and solve for x to find an x-intercept of (12, 0). Plot the two points and draw a line passing through the two points.

jimthompson5910 (jim_thompson5910):

so you had the coordinates mixed up

OpenStudy (anonymous):

ooh okay thanks again

jimthompson5910 (jim_thompson5910):

sure thing